A farmer grows a bushel of wheat and sells it to a miller for birr100. The miller turns the wheat into flour and then sells the flour to a baker for birr 300. The baker uses the flour to make bread and sells the bread to Rift Valley University for birr 600.The Rift Valley University students eat the bread. What is the value added by students? What is GDP?


1
Expert's answer
2022-01-13T09:00:05-0500

No value addition by the students as they just eat the bread and further processing of the product.


GDP= Final product cost=birr 600 or


Amount of value added by each person= (100- 0)+(300-100)+(600-300)


=100+200+300= birr 600